# Breeze-7B Technical Report

### v1

@@ 2.1 Model architecture customization
To increase the compression rate of the Traditional Chinese corpus, we employed a modified tokenizer with additional tokens, which is a strategy also present in domain adapted LLMs (Liu et al., 2023). This change has the extra benefit of speeding up both Traditional Chinese training and inference. We use the BPE algorithm (Sennrich et al., 2016) on a Chinese corpus to generate tokens to appended to the original tokenizer. After de-duplicating identical tokens already present in Mistral tokenizer, we extended 29873 tokens to it to arrive at 61872 total tokens. The input embedding layer and output embedding layer shapes are adjusted accordingly to accommodate the embeddings required for the new tokens. With the extended vocabulary, the compression rate on the Chinese corpus is around 2x compared to that of the Mistral model. As a result, our model enjoys a 11.1k context length on Traditional Chinese text, equivalent to around 10 pages of text, and we also observed a 2x speed up in the model training and inference speed.
@@ 4 Results
In this section, we reported the evaluated results of our model, and contrast our results with those of several notable contemporary models. Since some of the models have multiple variants, we select model variants that maximizes performance. For the language comprehension benchmarks, we use the instruct variants whenever possible; otherwise, we resort to the chat variants. We only use the Chat variants for the chatbot-oriented benchmarks. Additionally, to understand the effectiveness of the pretraining of different models, we also evaluate the language comprehension benchmarks on the base model. To increase robustness, we used a few-shot template to demonstrate the task format.
@@ 4.2.1 Traditional Chinese Language Comprehension benchmarks
Table 3: Benchmark results of pretrained models on TMMLU+ (5 shot) with Table (5 shot).
@@ Appendix A Chat Benchmark Details
Our GPT-4 judge incorporates the following prompt for MT-Bench-tw, where words in curly brackets will be replaced with real questions in the benchmark and answers of the language model. [[[[Instruction]]]]
@@
[[[[The End of Assistant’s Answer]]]]
Table  5 shows scores of the finetuned model on all 8 subcategories of the MT-Bench-tw benchmark. Breeze-7B-Instruct performs competitively with other models overall in the 7B size range, but it trails behind other models to greater extent on STEM and Writing subtasks. Enhancements in these areas is an exciting direction of future work.

### v2

@@ 2.1 Model architecture customization
To increase the compression rate of the Traditional Chinese corpus, we employed a modified tokenizer with additional tokens, which is a strategy also present in domain adapted LLMs (Liu et al., 2023). This change has the extra benefit of speeding up both Traditional Chinese training and inference. We use the BPE algorithm (Sennrich et al., 2016) on a Chinese corpus to generate tokens to appended to the original tokenizer. After de-duplicating identical tokens already present in Mistral tokenizer, we extended 29873 tokens to it to arrive at 61872 total tokens. The input embedding layer and output embedding layer shapes are adjusted accordingly to accommodate the embeddings required for the new tokens. With the extended vocabulary, the compression rate on the Chinese corpus is around 2x compared to that of the Mistral model. As a result, our model enjoys a 11.1k context length on Traditional Chinese text, equivalent to around 10 pages of text, and we also observed a 2x speed up in the model training and inference speed. In addition, our long model further quadruples the context length.
@@ 2.2 Training
2.3 Long Context Pretraining
The extension from short-context to long-context models is an important field of research, as this two-step approach has significant efficiency advantages over direct long-context training without performance degradation (Xiong et al., 2023). For this setting, different scaling methods exist (Chen et al., 2023; bloc, 97; Peng et al., 2023) to either induce a better inductive bias from the short-range trained positions or obtain better extrapolation properties (Liu et al., 2024). For our training, we incorporate a base-changing method, driven by the effectiveness and conciseness of the method on other open-sourced models (Jiang et al., 2023; AI et al., 2024). We trained the model on an additional 25 gigabytes of data consisting of Chinese, English and code data, and we upsampled long data for better long-range attention modeling. We turned off the sliding window function and set the context length to 32k, and keep all hyperparameters same as section 2.2
@@ 3.2 Chatbot-oriented benchmarks
3.3 Long-context benchmarks
To understand the models’ ability to attend to long context, we collected benchmarks that have inputs spanning over 10k tokens. At this scale, the benchmarks serve as a precursor to the performance on document-level tasks, as well as the commercially popular RAG (Retrieval Augmented Generation) use case. We first benchmark the model on the passkey retrieval task following previous work, and then on DRCD-long that is synthetically created from the Chinese benchmark DRCD. Passkey retrieval is a synthetic task to test whether a model could retrieve a 5-letter key within a long paragraph of irrelevant text. In addition to the variation of total sequence length, we additionally control the key position following observations that the depth of the key can significantly affect performance. DRCD-long extends the DRCD context to up to 64k Chinese characters with irrelevant articles, simulating inaccuracies in an RAG system. The task objective and scoring criterion remain the same as the original DRCD. We leave the curation details of this benchmark in Appendix B.
@@ 4 Results
In this section, we reported the evaluated results of our model, and contrasted our results with those of several notable contemporary models. Since some of the models have multiple variants, we select model variants that maximize performance. For the language comprehension benchmarks, we use the instruct variants whenever possible; otherwise, we resort to the chat variants. We only use the Chat variants for the chatbot-oriented benchmarks. Additionally, to understand the effectiveness of the pretraining of different models, we also evaluate the language comprehension benchmarks on the base model. To increase robustness, we used a few-shot template to demonstrate the task format.
@@ 4.2.1 Traditional Chinese Language Comprehension benchmarks
Table 3: Benchmark results of pretrained models on TMMLU+ (5 shot), DRCD (3 shot), Table (5 shot), and MMLU (5 shot).
@@ 4.2.2 Traditional Chinese Chatbot-oriented benchmarks
4.2.3 Long Context Benchmarks
We compare Breeze-7B-32k-Base with Breeze-7B-Base to show the effect after long context pre-training. As our focus of this paper is on enabling the long-context abilities of Breeze-7B-Base, we leave comparisons to other models as future work.
For the passkey retrieval task, we evenly split the key positions into 16 bins and tested 20 examples per bin. Accuracy is reported for each length-depth combination, and we visualize the results in Figure 2. The non-extended model can only consistently retrieve keys where they are located within a 4k token distance from the question, which is positioned at the end of the sequence. This phenomenon can be attributed to the sliding window size of non-extended model (4096), which limits the direct attention span, that is, information beyond this would have to rely on ”information forward passing”(Jiang et al., 2023) to be attended to. As expected, after long-context training, Breeze-7B-32k-Base performs consistently on the task up to 32k. An analysis of the incorrect examples reveals that almost all errors are copying errors, such as mistaking lowercase letters as uppercase, which indicates that the model has near-perfect performance in attending to the correct position.
Figure 2: Passkey Retrieval results of Breeze-7B-Base and Breeze-7B-32k-Base. The y-axis denotes the input sequence length, while the x-axis denotes the depth of the key position in the example. Each length-depth combination is trialed 20 times and the accuracy is color-coded with the colormap at the bottom.
For Long-DRCD, We observed in Table 5 that Breeze-7B-32k-Base has a large boost in performance in 3-shot DRCD compared to Breeze-7B-Base, and the model tested on DRCD-32k retains 77% of the DRCD performance. The ability to attend to text very far away benefits twofold: finding information that occurs in the front of the context and understanding the extractive setting of the task demonstrated by few-shot examples in the very front. We also include DRCD-64k which exceeds our training context length. At this length, the degradation is observable for both models, showing that our training method does not extrapolate well to longer sequences.
Model/Perforamce(EM) DRCD DRCD-16k DRCD-32k DRCD-64k
Breeze-7B-32k-Base-v1_0 79.73 69.68 61.55 25.82
Breeze-7B-Base-v1_0 80.61 21.79 15.29 12.63
Table 5: Benchmark results of pretrained models on DRCD with different context length (3-shot).
@@ Appendix A Chat Benchmark Details
Our GPT-4 judge incorporates the following prompt for MT-Bench-tw, where words in curly brackets will be replaced with real questions in the benchmark and answers of the language model. {mdframed} [[[[Instruction]]]]
@@
[[[[The End of Assistant’s Answer]]]] Table  6 shows scores of the finetuned model on all 8 subcategories of the MT-Bench-tw benchmark. Breeze-7B-Instruct performs competitively with other models overall in the 7B size range, but it trails behind other models to greater extent on STEM and Writing subtasks. Enhancements in these areas is an exciting direction of future work.
@@
Appendix B Long Context Benchmark Details
As seen in the passkey retrieval results, the position of the relevant information can greatly influence the performance. To ensure that the correct context span is even distributed with various depth across samples, our curation method of the DRCD-long tasks are as follow: For questions ${a,b,c,d,e...}$ that require context ${A,B,C,D,E...}$, we concatenate the contexts until they reach the target length to arrive at ${ABCDE,FGHIJ,...}$ as the contexts. The questions are automatically paired with the extended context. We have released DRCD-long at our huggingface page